OpenStudy (anonymous):

SOLVE SYSTEMS OF EQUATIONS BY USING SUBSTITUTION METHOED: -X+Y=5 Y=-3X+1 HELPPP

OpenStudy (radar):

With this on the second equation has y isolated so substitute for y in the first equation. -X +(-3X+1) = 5 -X -3X +1 =5 -4X = 5 - 1 -4x = 4 x=-1 Now substitute the value of X into the first equation -(-1) + Y = 5 1 + Y = 5 Y =5 -1= 4 Y = 4

OpenStudy (anonymous):

thank you soooo much